### v1

@@ 6 Numerical experiments
This section describes four numerical experiments with the variational EM algorithm. The first experiment (Sec. 6.1) uses simulated Markov chain data. The middle two experiments (Secs. 6.2 and  6.3) use real-world observations. The last experiment (Sec. 6.4) uses simulated genetics data. Code to run the experiments is available at https://github.com/chris-miles/MarkovChain-VEM.
6.1 Simulated Markov chain data
@@
$\operatorname{acc}=\frac{1}{N}\sum\nolimits_{n=1}^{N}\mathds{1}\{Z^{n}=\hat{Z}^{n}\},$ (7)
@@
The next experiment uses a publicly available data set that records the listening histories of Last.fm users in 2007. The goal of this experiment is to compare the accuracy of variational EM against previously published Markov chain mixture modeling results [37, 36] that also analyzed the Last.fm data.
Following the previous work [36], the ground-truth model is specified as follows. There are $k_{\mathrm{true}}=10$ Markov chains, corresponding to the Last.fm users with the greatest number of song listens. There are $s=100$ state possibilities, based on assigning a discrete label to each song in a user’s listening history by the predominant genre. If a user sequentially listens to songs from the same genre, these songs are collapsed into a single state. Hence, each Markov chain models the genre transitions for a Last.fm user. To make the learning problem difficult, each user’s listening history is broken into $75$ equal-length segments, and the segments are truncated to a short trajectory length $20\leq T\leq 100$.
Figure 5: Classification accuracy for the top 10 Last.fm users, given $N=75$ trajectories and $s=100$ possible genres for each user. Top: Variational EM identifies is twice as accurate as the best method reported in [36]. Bottom: Confusion matrix for identifying the top 10 users with $T=250$.
The results in Fig. 5 show that variational EM with $k=k_{\mathrm{true}}=10$ components is twice as accurate as the best results reported in [36]. Variational EM is also more accurate than the work [37], which reports $90\%$ misclassifications on the Las.fm data set with $k_{\mathrm{true}}=10$ components. The comparison with [37] is slightly complicated however and is not reported in the figure, since the trajectories in [37] are more numerous and shorter ($N=30,000$, $T=3$).
@@
For further insight, a new data set was prepared with a length $T=1000$ trajectory for each of the top $N=497$ Last.fm users. Variational EM was applied to this data set with a maximum number of $k=100$ components. Fig. 6 shows the mixture modeling results, which indicates that a few components have substantially large memberships. The 3 largest components correspond to well-known genres of ‘indie’, ‘electronic’, and ‘metal’. The overwhelming popularity of the indie component explains why distinguishing Last.fm users may be difficult. Many indie rock listeners had quite similar listening histories in 2007 [50].
@@ 6.3 Ultrarunners data set
The third experiment is based on data from the 2012 International Association of Ultrarunners (IAU) World Championship held in Katowice, Poland [51, 52]. The goal of this experiment is to determine how many pacing patterns can be observed in the data and which pacing patterns lead to the best overall performance. Conventional wisdom among ultrarunners contends that a slower pace to start is ideal [53], as many runners tend to overexert themselves during the start and tire later during a race.
Pacing is an individualized measure of performance, so each runner’s data over the 24-hour event was normalized based on their average speed. First, $12$ runners that did not complete a single lap were removed from the data set, yielding $N=248$ total trajectories. Next, each runner’s average speed was calculated by dividing the total number of laps over the 2.314km course by the total number of hours until finishing or dropping out of the event (typically 24 hours). Individualized speeds were calculated for each runner by dividing the number of laps completed each hour by the average speed. Last, the data was categorized into $s=4$ Markov states: “resting” when the hourly speed is less than $0.5$ times the average speed, “normal” when the hourly speed is between $0.5$ and $1.5$ times the average speed, “strained” when the hourly speed is greater than $1.5$ times the average speed, and a terminal “ended” state.
@@
The variational EM algorithm identifies three distinct running patterns in the data, as displayed in Fig.  7. The predominant pattern ($70\%$ of runners) involves a roughly constant pace, with slightly higher relative speed at the start of the race. The second-most-common pattern ($25\%$ of runners) involves overexertion at the start of the race and a significant slowdown later. The third pattern involves erratic running with no apparent strategy. A previous analysis identified three similar running patterns in the 2012 ultrarunner data [51].
@@
6.4 Simulated gene expression data
@@
The last experiment is based on simulations of the mutual-inhibition, self-activation (MISA) gene circuit. The circuit consists of an $A$ gene and a $B$ gene that inhibit each other and activate themselves through the production of $a$ and $b$ proteins. The MISA circuit has been observed in many biological systems [54, 55, 56, 57] and studied extensively by theorists [58, 59, 60, 61]. The goal of the experiment is to test whether variational EM can correctly classify data from this biologically important system. The experiment also tests whether spectral clustering (Alg. 1) can automatically identify Markov states from the MISA data.
The MISA circuit is typically described as a chemical reaction network [62] containing the following species and reactions. There is one $A$ gene and one $B$ gene that exhibit a range of conditions $ij\in\{00,01,10,11\}$, in which they produce proteins at a rate $g_{ij}$:
$A_{ij}$ $\overset{g_{ij}}{\to}A_{ij}+a,\quad ij={00,01,01,11},$
$B_{ij}$ $\overset{g_{ij}}{\to}B_{ij}+b,\quad ij={00,01,01,11}.$
The rates are fixed at $g_{00}=g_{01}=g_{11}=10$ and $g_{10}=100$ since $i=1$ indicates the presence of an activator and $j=1$ indicates the presence of a repressor. The proteins degrade at rate $d=1$:
[…]
Last, as the proteins are produced and degrade, they influence the conditions of the $A$ and $B$ genes as follows:
$A_{0j}+2a$ $\stackrel{{\scriptstyle h_{a}}}{{\underset{f_{a}}{\rightleftarrows}}}A_{1j},\quad B_{0j}+2b\stackrel{{\scriptstyle h_{a}}}{{\underset{f_{a}}{\rightleftarrows}}}B_{1j},\quad j=0,1.$
@@
Three of the rate parameters are fixed to $h_{a}=10^{-1}$, $f_{a}=1$, and $h_{r}=10^{-3}$. However, $f_{r}$ is a free parameter that controls the extent of protein activation versus repression in the system. In summary, the MISA gene circuit can be encoded as a vector $X_{t}\in\{00,01,10,11\}^{2}\times\mathbb{N}^{2}$, where the first two coordinates indicate conditions of the $A$ and $B$ genes and the last two coordinates indicate populations of $a$ and $b$ proteins. To generate MISA trajectories, the system $X_{t}$ was run forward in time with various values of the $f_{r}$ parameter using the stochastic simulation algorithm (SSA) implemented in PyGillespie [63]. The algorithm produces a continuous-time trajectory $(X_{t})_{t\geq 0}$; however, the states were sampled at uniformly spaced times $t=0,1,2,\ldots$ to produce a discrete-time data set.
@@
The MISA gene circuit is interesting to biologists because of its emergent stochastic-switching behavior, which is useful in modeling cell-fate decisions [5]. Fig. 8 displays the stochastic switching in the populations of $a$ and $b$ proteins. The protein populations toggle between four metastable states, which are associated with each gene being “on” or “off”. When the rate $f_{r}$ is low (top panels), the genes are frequently off, leading to protein populations of just 0–20. When $f_{r}$ is high (bottom panels), both genes are frequently on, leading to increased populations of 50–150 proteins. The most interesting behavior occurs for intermediate values of $f_{r}$ (middle panels), because then one gene is typically on and the other is off. The genes actively compete to produce more proteins.
@@
The MISA gene circuit is not Markovian in the protein numbers $a$ and $b$. Nonetheless, a Markov state model can be constructed by applying spectral clustering (Alg. 1) to the $(a,b)$ data. The middle panel of Fig. 8 shows the outcome of spectral clustering using a Gaussian kernel with bandwidth $\sigma=50$. The four identified clusters are interpretable and they efficiently represent the metastable dynamics.
@@
Figure 9: Classification accuracy for simulated MISA data containing 15 trajectories with $f_{r}^{(1)}=0.01$ and 15 trajectories with $f_{r}^{(2)}\neq f_{r}^{(1)}$. Trajectories are reliably identified when $T\geq 25$ and $f_{r}^{(2)}/f_{r}^{(1)}\geq 25$.
@@
Variational EM can be used to untangle multiple Markov chains from a single MISA data set. This application area is increasingly important, as MISA models with variable parameters are now being engineered in labs [64] and applied to understand long chains of cell differentiation [56]. As an example application, researchers have obtained time series data for a MISA gene circuit in the bacteriophage $\lambda$ switch [8]. They analyzed the time-series data assuming a single Markov chain, but the Markov chain mixture model provides the ability to discern heterogeneities within the samples.
@@
To test the feasibility of the mixture modeling approach, spectral clustering and variational EM were applied to a range of simulated MISA data sets containing $15$ trajectories with an unbinding rate $f_{r}^{(1)}=0.01$ and $15$ trajectories with a different unbinding rate $f_{r}^{(2)}\neq f_{r}^{(1)}$. The algorithms were used to untangle the two populations of trajectories, leading to the results in Fig. 9. The results indicate that the classification accuracy depends greatly on the trajectory length $T$ and the parameter ratio $f_{r}^{(2)}/f_{r}^{(1)}$. When $f_{r}^{(2)}/f_{r}^{(1)}\geq 25$ and $T\geq 25$, the populations can be separated with near-perfect accuracy. However, short trajectories with $T\leq 5$ cannot be reliably separated for any $f_{r}^{(2)}$ parameter. These results support the theoretical analysis in Thm. 1 that links the classification accuracy with the Kullback-Leibler divergence, which increases with $f_{r}^{(2)}/f_{r}^{(1)}$ and increases exponentially fast with $T$.
@@ 7 Conclusion
This paper has proposed an extension of Markov state modeling that enables the study of heterogeneities within time series data. Previous work employed Markov state modeling under the assumption that the observations come from a homogeneous population, modeled by a single Markov chain. This paper establishes the theory and practice of learning a mixture of different Markov chains simultaneously.
The paper has motivated and tested a variational EM algorithm that automatically identifies the number of chains and the dynamics of each chain. The algorithm is computationally efficient: unlike past work, it identifies the number of chains organically, without relying on expensive model comparisons or posterior sampling.
@@
The paper also contributes a new theorem that fundamentally limits the classification accuracy of any Markov chain classification algorithm. The bound is stated in terms of the Kullback-Leibler divergence between the underlying Markov chains, and it suggests exponentially increasing accuracy with the trajectory length $T$. This prediction is repeatedly supported in the numerical experiments (Secs. 6.1, 6.2, 6.4). The theory precisely quantifies the established wisdom [30] that long trajectories are better than short ones.
Last, there is the question of how the Markov chain mixtures can be used by scientists. Three of the four experiments (Secs. 6.2, 6.3, 6.4) were designed with applications in mind, and the method was found to yield sensible, reliable interpretations. One of the authors (C.E.M.) listened to indie rock on Last.fm in 2007 while the other author (R.J.W.) runs ultramarathons competitively, and they can attest to the qualitative accuracy of the Markov chain mixture models.
In summary, Markov chain mixture models have an established track record in distinguishing heterogeneities in human behavior, such as surfing the internet [22], commuting between home, work, and school [27], listening to music (Sec. 6.2), and running ultramarathons (Sec. 6.3). The variational EM algorithm is the best currently available strategy for fitting these models efficiently and correctly. Moving forward, there is the potential for broad utility of Markov chain mixture model in other areas (chemistry, biology, climate science) where Markov state modeling is currently popular. Indeed, the experiments in Sec. 6.4 demonstrate promising first steps toward applications in the rapidly evolving area of gene expression data analysis [65, 66, 67, 68].
@@ Acknowledgements
The authors thank Elizabeth Read and Adam MacLean for helpful discussions about gene expression time series. C.E.M. was partially supported by a University of California Society of Hellman Fellows fund. R.J.W. was supported by the Office of Naval Research through BRC Award N00014-18-1-2363, the National Science Foundation through FRG Award 1952777, and Caltech through the Carver Mead New Adventures Fund, under the aegis of Joel A. Tropp.

### v2

@@ 5 Numerical Experiments
This section describes four numerical experiments with the variational EM algorithm. The first experiment (Sec. 5.1) uses synthetic Markov chain data. The middle two experiments (Secs. 5.2 and  5.3) use real-world observations. The last experiment (Sec. 5.4) uses synthetic genetics data. Code to run the experiments is available at https://github.com/chris-miles/MarkovChain-VEM.
5.1 Synthetic Markov Chain Data
@@
$\operatorname{acc}=\frac{1}{N}\sum_{n=1}^{N}\mathds{1}\{Z^{n}=\hat{Z}^{n}\},$ (7)
@@
The next experiment uses a publicly available data set that records the listening histories of Last.fm users in 2007. The works [34, 35] report trajectory-level clustering results on this Last.fm Markov-chain-mixture benchmark, making them natural baselines.
Following the previous work [34], the ground-truth model is specified as follows. The data contains $k_{\mathrm{true}}=10$ Markov chains, corresponding to the Last.fm users with the greatest number of song listens. The data is coded using $s=100$ states based on assigning a discrete label to each song in a user’s listening history by the predominant genre. If a user sequentially listens to songs from the same genre, these songs are collapsed into a single state. Hence, each Markov chain models the genre transitions for a Last.fm user. To make the learning problem difficult, each user’s listening history is broken into $75$ equal-length segments, and the segments are truncated to a short trajectory length $20\leq T\leq 100$.
Figure 5: Classification accuracy for the top 10 Last.fm users, given $N=75$ trajectories and $s=100$ possible genres for each user. Top: Variational EM is twice as accurate as the best method reported in [34]. Bottom: Confusion matrix for identifying the top 10 users with $T=250$.
Fig. 5 shows that the Bayesian variational EM approach with $k=k_{\mathrm{true}}=10$ components achieves classification accuracy approximately $2\times$ higher than the experiments in [34]. The error is also much lower than the $90\%$ classification error in [35]. These results are included for context, although the earliest benchmark [35] used more numerous and shorter trajectories ($N=30,000$, $T=3$), so the numbers are not directly comparable.
@@
For further insight, a new data set was prepared with a length $T=1000$ trajectory for each of the top $N=497$ Last.fm users. Variational EM was applied to this data set with a maximum number of $k=100$ components. Fig. 6 shows the mixture modeling results, which indicate that a few components have substantially large memberships. The 3 largest components correspond to well-known genres of ‘indie’, ‘electronic’, and ‘metal’. The overwhelming popularity of the indie component explains why distinguishing Last.fm users may be difficult. Many indie rock listeners had quite similar listening histories in 2007 [49].
@@ 5.3 Ultrarunners data set
The third experiment is based on data from the 2012 International Association of Ultrarunners (IAU) World Championship held in Katowice, Poland [50, 51]. The goal of this experiment is to find the predominant pacing patterns in the data and infer which pacing patterns lead to the best overall performance. Conventional wisdom among ultrarunners contends that a slower pace to start is ideal [52], as many runners tend to overexert themselves during the start and tire later during a race.
Pacing is an individualized measure of performance, so each runner’s data over the 24-hour event was normalized based on average speed. First, $12$ runners that did not complete a single lap were removed from the data set, yielding $N=248$ total trajectories. Then, each runner’s average speed was calculated by dividing the total number of laps over the 2.314 km course by the total number of hours until finishing or dropping out of the event (typically 24 hours). Each runner’s data was normalized by dividing the number of laps per hour by the average speed. Last, the data was categorized into $s=4$ Markov states: “resting” when the hourly speed is less than $0.5$ times the average speed, “normal” when the hourly speed is between $0.5$ and $1.5$ times the average speed, “strained” when the hourly speed is greater than $1.5$ times the average speed, and a terminal “ended” state.
@@
The variational EM algorithm identifies three distinct running patterns in the data, as displayed in Fig.  7. The predominant pattern ($70\%$ of runners) involves a roughly constant pace, with slightly higher relative speed at the start of the race. The second-most-common pattern ($25\%$ of runners) involves overexertion at the start of the race and a significant slowdown later. The third pattern involves erratic running with no apparent strategy. A previous analysis identified similar running patterns in the 2012 ultrarunner data [50].
@@
5.4 Synthetic Gene Expression Data
@@
The last experiment is based on synthetic data generated from the mutual-inhibition, self-activation (MISA) gene circuit. The MISA circuit has been observed in many biological systems [53, 54, 55, 56], and it has been studied extensively by theorists [57, 58, 59, 60]. It consists of an $A$ gene and a $B$ gene that inhibit each other and activate themselves through the production of $a$ and $b$ proteins. Mathematically, it can be modeled as a chemical reaction network [61] containing the following species and reactions. There is one $A$ gene and one $B$ gene that exhibit conditions $ij\in\{00,01,10,11\}$, where $i=1$ indicates the presence of an activator and $j=1$ indicates the presence of a repressor. The A and B genes produce proteins at a rate $g_{ij}$:
$A_{ij}$ $\overset{g_{ij}}{\to}A_{ij}+a,\quad ij={00,01,10,11},$
$B_{ij}$ $\overset{g_{ij}}{\to}B_{ij}+b,\quad ij={00,01,10,11},$
where $g_{00}=g_{01}=g_{11}=10$ and $g_{10}=100$. The proteins degrade at rate $d=1$:
$a\overset{d}{\to}\varnothing,\quad b\overset{d}{\to}\varnothing.$
Last, the proteins influence the conditions of the $A$ gene and $B$ gene as follows:
$A_{0j}+2a$ $\stackrel{{\scriptstyle h_{a}}}{{\underset{f_{a}}{\rightleftarrows}}}A_{1j},\quad B_{0j}+2b\stackrel{{\scriptstyle h_{a}}}{{\underset{f_{a}}{\rightleftarrows}}}B_{1j},\quad j=0,1,$
@@
Three of the rate parameters are fixed to $h_{a}=10^{-1}$, $f_{a}=1$, and $h_{r}=10^{-3}$. However, $f_{r}$ is a free parameter that controls the extent of protein activation versus repression in the system. In summary, the MISA gene circuit can be encoded as a vector $X_{t}\in\{00,01,10,11\}^{2}\times\mathbb{N}^{2}$. The first two coordinates indicate the conditions of the $A$ gene and $B$ gene, and the last two coordinates indicate the populations of the $a$ and $b$ proteins.
The MISA trajectories were simulated using the stochastic simulation algorithm (SSA) implemented in PyGillespie [62] with various values of the $f_{r}$ parameter. The SSA algorithm produces a continuous-time trajectory $(X_{t})_{t\geq 0}$; however, the states were sampled at uniformly spaced times $t=0,1,2,\ldots$ to produce a discrete-time data set.
@@
Fig. 8 displays the stochastic switching in the populations of $a$ and $b$ proteins. The protein populations toggle between four metastable states, which are associated with each gene being “on” or “off”. When the rate $f_{r}$ is low (top panels), the genes are frequently off, leading to protein populations of just 0–20. When $f_{r}$ is high (bottom panels), both genes are frequently on, leading to increased populations of 50–150 proteins. The most interesting behavior occurs for intermediate values of $f_{r}$ (middle panels), because then one gene is typically on and the other is off. The genes actively compete to produce more proteins.
@@
The MISA gene circuit is not Markovian in the protein numbers $a$ and $b$. Nonetheless, a Markov state model can be constructed by applying spectral clustering [28] to the $(a,b)$ data. The middle panel of Fig. 8 shows the outcome of spectral clustering using a Gaussian kernel with bandwidth $\sigma=50$. The four identified clusters are interpretable and they efficiently represent the metastable dynamics.
@@
Figure 9: Classification accuracy for synthetic MISA data containing 15 trajectories with $f_{r}^{(1)}=0.01$ and 15 trajectories with $f_{r}^{(2)}\neq f_{r}^{(1)}$. Trajectories are reliably identified when $T\geq 25$ and $f_{r}^{(2)}/f_{r}^{(1)}\geq 25$.
@@
Variational EM can be used to untangle multiple Markov chains from a single MISA data set. This application area is increasingly important, as MISA models with variable parameters are now being engineered in labs [63] and applied to understand long chains of cell differentiation [55]. As an example application, researchers have obtained time-series data for a MISA gene circuit in the bacteriophage $\lambda$ switch [8]. They analyzed the time-series data assuming a single Markov chain, but the Markov chain mixture model provides the ability to discern heterogeneities within the samples.
@@
To test the feasibility of the mixture modeling approach, spectral clustering and variational EM were applied to a range of synthetic MISA data sets containing $15$ trajectories with an unbinding rate $f_{r}^{(1)}=0.01$ and $15$ trajectories with a different unbinding rate $f_{r}^{(2)}\neq f_{r}^{(1)}$. The algorithms were used to untangle the two populations of trajectories, leading to the results in Fig. 9. The results indicate that the classification accuracy depends greatly on the trajectory length $T$ and the parameter ratio $f_{r}^{(2)}/f_{r}^{(1)}$. When $f_{r}^{(2)}/f_{r}^{(1)}\geq 25$ and $T\geq 25$, the populations can be separated with near-perfect accuracy. However, short trajectories with $T\leq 5$ cannot be reliably separated for any $f_{r}^{(2)}$ parameter. These results support the theoretical analysis in Thm. 1, since the Kullback-Leibler divergence increases with $f_{r}^{(2)}/f_{r}^{(1)}$ and increases linearly with $T$.
@@ 6 Conclusion
This paper has proposed an extension of Markov state modeling that enables the study of heterogeneities within time-series data. Previous work employed Markov state modeling under the assumption that the observations come from a homogeneous population modeled by a single Markov chain. This paper develops the theory and practice of learning a mixture of different Markov chains simultaneously.
The paper has motivated and tested a variational EM algorithm that automatically identifies the number of chains and the dynamics of each chain. The proposed mixture modeling approach deliberately combines classical, well-understood components, based on the discretization step from Markov state modeling and the fitting step from variational EM. It contributes an interpretable and competitive baseline for heterogeneous dynamical data. The algorithm is computationally efficient: unlike past work, it identifies the number of chains organically, without relying on expensive model comparisons or posterior sampling.
@@
The paper also provides a theorem that lower bounds the classification error for any Markov chain mixture model, similar to classical information-theoretic bounds. The bound is stated in terms of the Kullback-Leibler divergence between the underlying Markov chains, and it suggests the optimal classification error decreases exponentially in $T$ because the trajectory-level KL divergence grows linearly in $T$. This prediction is repeatedly supported in the numerical experiments (Secs. 5.1, 5.2, 5.4). The theory precisely quantifies the established wisdom [13] that long trajectories are better than short ones. The bound requires no ergodicity or mixing-time assumptions, making it applicable to transient or non-stationary dynamics common in short experimental trajectories.
Last, there is a natural question: when should scientists use Markov chain mixture models in the future? Three of the four experiments (Secs. 5.2, 5.3, 5.4) involved real data or real biological systems, and the mixture modeling approach led to accurate and interpretable results. One of the authors (C.E.M.) listened to indie rock on Last.fm in 2007 while the other author (R.J.W.) runs ultramarathons competitively, and they can attest to the qualitative accuracy of the Markov chain mixture models. More broadly, these results reinforce a simple point: careful probabilistic modeling and inference, even when built from classical ingredients, can remain competitive in modern high-throughput settings, serving as an interpretable baseline against which newer approaches are compared.
In conclusion, Markov chain mixture models have a track record of distinguishing meaningful heterogeneities in human behavior, for example, heterogeneous patterns of surfing the internet [15], commuting between home, work, and school [16], listening to music (Sec. 5.2), or running ultramarathons (Sec. 5.3). The variational EM algorithm is an efficient strategy for fitting these models while automatically selecting the number of components. Moving forward, Markov chain mixture models could be extended to other areas (chemistry, biology, climate science) where Markov state modeling is already popular. Indeed, the experiments in Sec. 5.4 demonstrate promising first steps toward applications in the rapidly evolving area of gene expression data analysis [64, 65, 66, 67].
@@ Acknowledgements
The authors thank Elizabeth Read and Adam MacLean for helpful discussions about gene expression time series. C.E.M. was partially supported by a University of California Society of Hellman Fellows fund and NSF CAREER DMS-2339241. R.J.W. was supported by the Office of Naval Research through BRC Award N00014-18-1-2363, the National Science Foundation through FRG Award 1952777, and Caltech through the Carver Mead New Adventures Fund, under the aegis of Joel A. Tropp.